OpenStudy (kmallette09):

At a basketball game, members of the Spirit Club use a t-shirt cannon to launch team t-shirts to spectators in the stands. The equation below describes the height, h, of the t-shirt, t seconds after it is shot from the cannon c feet above the ground, at an initial velocity of v. h = -16t 2 + vt + c Solve the equation for initial velocity, v. The cannon launched a t-shirt from an initial height of 5 feet. After 3 seconds the shirt has reached a height of 131 feet. Use your equation from #1 to find the initial velocity of the t-shirt. Show your work. What does the equation from #1 tell y

OpenStudy (kmallette09):

tell you about the value of the velocity, v, when t = 0? Explain.

OpenStudy (kmallette09):

is that it

OpenStudy (will.h):

here's what i think h = -16t 2 + vt + c where h= 131 and t= 3 and C=5 now solve for v 131= -16(3)^2 + v(3) + 5 131= -144 + 3v + 5 131 + 144 -5 = 3v 270=3v 90=v Therefore it took a velocity of 90 and time of 3sec to reach 131 height.

OpenStudy (will.h):

that's the right one. :) sorry for being late

OpenStudy (will.h):

Thank you for the medal :)

OpenStudy (will.h):

post your last question before i go to sleep.

OpenStudy (kmallette09):

ok so is that the answer u think

OpenStudy (will.h):

tell you about the value of the velocity, v, when t = 0? Explain. you can say when they time is zero then h = -16t 2 + vt + c 131= -16(0)^2 + v(0) + 5 therefore 131= 0 + 0v + 5 so basically the velocity is zero when the time is zero

OpenStudy (will.h):

make sense?

OpenStudy (kmallette09):

yeah

OpenStudy (retireed):

I might be confused... because I usually am, but 90 is the velocity at 3 secs into the flight, so the initial velocity should be greater? You need to differentiate the distance equation to find the velocity at t = 0